# Efficiently Identifying Watermarked Segments in Mixed-Source Texts

### v1

@@ Abstract
Text watermarks in large language models (LLMs) are increasingly used to detect synthetic text, mitigating misuse cases like fake news and academic dishonesty. While existing watermarking detection techniques primarily focus on classifying entire documents as watermarked or not, they often neglect the common scenario of identifying individual watermark segments within longer, mixed-source documents. Drawing inspiration from plagiarism detection systems, we propose two novel methods for partial watermark detection. First, we develop a geometry cover detection framework aimed at determining whether there is a watermark segment in long text. Second, we introduce an adaptive online learning algorithm to pinpoint the precise location of watermark segments within the text. Evaluated on three popular watermarking techniques (KGW-Watermark, Unigram-Watermark, and Gumbel-Watermark), our approach achieves high accuracy, significantly outperforming baseline methods. Moreover, our framework is adaptable to other watermarking techniques, offering new insights for precise watermark detection.
@@ 1 Introduction
However, identifying AI-generated text is becoming increasingly difficult as LLMs reach human-like proficiency in various tasks. One line of research (OpenAI, 2023; Tian, 2023; Mitchell et al., 2023) trains machine learning models as AI detectors by collecting datasets consisting of both human and LLM-generated texts. Unfortunately, these approaches are often fragile (Shi et al., 2024) and error-prone (Liang et al., 2023), ultimately leading OpenAI to terminate its deployed detector (Kelly, 2023). Watermarking has emerged as a promising solution to this challenge. By embedding identifiable patterns or markers within the generated text, watermarks can signal whether a piece of text originates from an LLM.
Existing watermark detection methods (Aaronson, 2023; Kirchenbauer et al., 2023; Zhao et al., 2023; Kuditipudi et al., 2023; Christ et al., 2023; Hu et al., 2024) are primarily designed for text-level classification, labeling a piece of text as either watermarked or not. However, these methods are insufficient for many real-world scenarios where documents contain mixed-source texts, and only specific sections are LLM-generated. For instance, malicious actors might use LLMs to manipulate certain sections of a news article to spread misinformation. Detecting watermarks within long, mixed-source texts presents a significant challenge, especially when aiming for subsequence-level detection with uncertainty quantification, similar to plagiarism detection systems like “Turnitin [^1] https://www.turnitin.com”. This is because the watermarked signal may be weakened throughout the increasing text length and may not be easily identifiable using conventional detection methods.
@@ 2.1 LLM Watermark and Detection
Red-Green (statistical) watermarking methods partition the vocabulary into two sets, “green” and “red”, using a pseudorandom function $R(h,{\sf k},\gamma)$. This function takes as input the length of the preceding token sequence ($h$), a secret watermark key (${\sf k}$), and the target proportion of green tokens ($\gamma$). During text generation, the logits of green tokens are subtly increased by a small value $\delta$, resulting in a higher proportion of green tokens in the watermarked text compared to non-watermarked text. Two prominent Red-Green watermarking methods are KGW-Watermark (Kirchenbauer et al., 2023, 2024) and Unigram-Watermark (Zhao et al., 2023). KGW-Watermark utilizes $h\geq 1$, considering the prefix for hashing. Unigram-Watermark employs fixed green and red lists, disregarding previous tokens by effectively setting $h=0$ to enhance robustness. Watermark detection in both methods involves identifying each token’s membership in the green or red list
@@ 3 Method
3.1 Problem Statement
Identifying watermark segments within a long text sequence $y$ presents two key challenges. First, we need to design a classification rule $\mathcal{M}(x)\rightarrow\{0,1\}$ that determines whether $y$ contains a watermark segment. To address this, we propose the Geometric Cover Detector (GCD), which enables multi-scale watermark detection. Second, accurately locating the watermark segments $y_{s_{i}:e_{i}}$ within the full sequence $y$ requires finding the start and end token indices, $s_{i}$ and $e_{i}$, for each watermark segment. We introduce the Adaptive Online Locator (AOL) with the Aligator algorithm to precisely identify the position of the watermarked text span within the longer sequence.
Figure 1: Illustration of the watermark segment detection process. The input sequence could be mixed-source of watermark text and unwatermark text. The input sequence could be a mixed-source of watermarked text and unwatermarked text. We use geometric covers to partition the text and detect watermarks in intervals. We also formulate localization as an online denoising problem to reduce computational complexity. The example shown is drawn from the abstract of Bengio et al. (2024), with the watermarked part generated by a watermarked Mistral-7B model.
Algorithm 1 Geometry Cover Watermark Detection
Input: Mixed-source text $y$ of length $n$, target false positive rate (FPR) $\tau$, watermark detector score function $\operatorname{Score}$, FPR calibration function $F$
@@ 3.2 Watermark Segment Classification
Algorithm 1 outlines our approach. For each segment $\mathcal{I}_{t}:y_{i_{t}:j_{t}}$ in the GC, we first compute a detection score using the appropriate watermark detector for the scheme employed (e.g., Equation 1 for Red-Green Watermark or Equation 2 for Gumbel Watermark). This score, along with the segment itself, is then passed to an FPR calibration function $F$. This function estimates the FPR associated with the segment. Further details on FPR calibration can be found in the Appendix A.1.
@@ 3.3 Precise Watermark Position Localization
Input: Mixed-source text $y$, threshold $\zeta$, iterations $m$
@@
Furthermore, relying solely on individual token scores for localization is unreliable due to the inherent noise in the watermarking process. To address this issue, we propose to formulate it as a sequence denoising problem (a.k.a., smoothing or nonparametric regression) so we can provide a pointwise estimate of the expected detection score for each token. Specifically, the denoising algorithm tasks a sequence of noisy observations $s_{1},...,s_{n}$ and output $\{\theta_{t}\}_{t\in[n]}$ as an estimate to $\{\mathop{\mathbb{E}}[s_{t}]\}_{t\in[n]}$.
@@
$\frac{1}{n}\sum_{t}(\theta_{t}-\mathop{\mathbb{E}}[s_{t}])^{2}=\tilde{O}\left(\min\left\{n^{-1}(1+\sum_{t=2}^{n}\mathbf{1}_{\mathop{\mathbb{E}}[s_{t}]\neq\mathop{\mathbb{E}}[s_{t-1}]}),n^{-1}\vee n^{-2/3}(\sum_{t=2}^{n}\left|\mathop{\mathbb{E}}[s_{t}]-\mathop{\mathbb{E}}[s_{t-1}]\right|)\right\}\right).$
@@
$\frac{1}{j-1}\sum_{t=i}^{j}(\theta_{t}-\frac{1}{j-i}\sum_{t^{\prime}=i}^{j}\mathop{\mathbb{E}}[s_{t^{\prime}}])^{2}\leq\tilde{O}(1/\sqrt{j-i}).$
@@ 4 Experiment
In this section, we discuss the detailed experiment settings and then show the detection results for watermark segment classification and precise watermark position localization. We also consider different lengths of the mixed-source text and show the detection results. Furthermore, we conduct adversarial attacks to try to remove the watermark and then show the detection robustness results.
4.1 Datasets and Mixed-source Texts
We utilize two text datasets: C4 (Raffel et al., 2020) and Arxiv (Cohan et al., 2018). The “Colossal Clean Crawled Corpus” (C4) dataset is a collection of English-language text sourced from the public Common Crawl web scrape, a rich source for unwatermarked human-written text. We use random samples from the news-like subset of the C4 dataset in our experiments. The Arxiv dataset is part of the Scientific-Papers dataset collected from scientific repositories, arXiv.org and PubMed.com. We use the Arxiv split in our experiments, which contains abstracts and articles of scientific papers. Both datasets are used to construct watermarked positive samples and human-written negative samples.
To transform unwatermarked samples into partially watermarked samples, we randomly select 3-5 sentences in a long text and set them as prompts. Then, we generate 300 tokens of watermarked text conditioned on the prompts using large language models. The generated responses replace the original suffix sentences after the prompt. In this way, we embed 300-token watermarks into 3000-token contexts from the datasets, making the watermark 10% of the mix-sourced text. We randomly choose the position of the watermark in this longer context and record the locations for later testing. Our goal is to determine if a document contains watermark text and locate its position. For each dataset, we use 500 samples as the test set to show the results.
4.2 Language Models and Watermarking Methods.
We use the publicly available LLaMA-7B (Touvron et al., 2023) and Mistral (Jiang et al., 2023) models. To verify the general applicability of the watermark detection methods, we select three watermarking techniques: Gumbel-Watermark (Aaronson, 2023), KGW-Watermark (Kirchenbauer et al., 2023), and Unigram-Watermark (Zhao et al., 2023). These methods represent the state-of-the-art watermarking approaches for large language models, offering high quality, detectability, and robustness against adversarial attacks. For all watermarking generations, we configure the temperature to 1.0 for multinomial sampling. Additionally, for KGW-Watermark and Unigram-Watermark, we set the green token ratio $\gamma$ to 0.5 and the perturbation $\delta$ to 2.0.
4.3 Baselines
In watermark segment detection, we use the original watermark detector in each watermarking method as the Vanilla baseline to compare with our approach GCD. In watermark segment localization, we use RoBERTa (Liu et al., 2019) models for comparing with our method AOL. We train each RoBERTa (designed for different watermarking methods) to predict whether a sequence is a watermarked sequence or not, given the watermark detection scores $r$ for each token. We add an extra fully connected layer after getting the representation of the [CLS] token. We construct 1000 training samples with 60 token scores as input and the binary label of this segment as the label. We train the RoBERTa model for 20 epochs and enable early stopping if the loss converges. It can reach over 90% accuracy in the training set. During testing on mixed-source text, we employ the sliding window idea to test each chunk for watermarks and then calculate the IoU score.
4.4 Evaluation
For the watermarked text classification task, we report the true positive rates (TPR) based on different specified false positive rates (FPR). Maintaining a low FPR is critical to ensure that human-written text is rarely misclassified as LLM-generated text.
Since the FPR at the per-instance level differs from the document-level FPR, we calibrate FPR to three distinct levels in each scenario to enable fair comparisons. Specifically, we manipulate the pre-segment FPR (Seg-FPR) by adjusting the threshold parameter $\tau$ as outlined in Algorithm 1. Then, we can get the empirical document FPR (Doc-FPR) by evaluating our method GCD based on pure natural text. For Vanilla, we set the FPR according to GCD’s empirical FPR and subsequently test for its empirical TPR.
For locating specific watermark segments, we calculate the Intersection over Union (IoU) score to measure the accuracy of watermark segment localization. The IoU score computes the ratio of the intersection and union between the ground truth and inference, serving as one of the main metrics for evaluating the accuracy of object detection algorithms:
$\operatorname{IoU}$ $=\frac{\text{Area of Intersection}}{\text{Area of Union}}=\frac{|\text{Detected Tokens}\cap\text{Watermarked Tokens}|}{|\text{Detected Tokens}\cup\text{Watermarked Tokens}|}$
4.5 Detection Results
4.5.1 Watermark Segment Classification Results
@@
4.5.2 Precise Watermark Position Localization Results
For the watermark position localization task, we evaluate our proposed method AOL against the baseline method RoBERTa (Table 2). We calculate the average IoU score to quantify the precision of the watermark localization.
Our method consistently outperforms the baseline across all test settings. For example, on the C4 dataset using the mistral-7B model, AOL achieves a substantially higher IoU score of 0.809 compared to 0.301 for RoBERTa. We also test AOL’s ability to detect multiple watermarks by inserting 3x300-token Gumbel watermarks (generated by Mistral-7B) into 6000-token texts. Across 200 samples, the average IoU for detecting the watermarks is 0.802, demonstrating AOL’s effectiveness for multiple watermark detection.
Figure 2 provides a case example illustrating the improved localization performance of AOL on the Gumbel watermark with the Mistral-7B model. The upper image shows the boundary effects of using online learning. The lower image demonstrates more precise localization resulting from the circular starting strategy with 10 random starting points.
@@ 4.5.2 Precise Watermark Position Localization Results
4.6 Detection Results with Different Lengths
@@
Table 3: Vanilla and GCD watermark segment classification result with Unigram Watermark on Mistral-7B with different target false positive rates.
@@ 4.7 Detection Robustness Against Attacks
Table 4: Watermark segment classification and localization performance with different attacks.
We evaluate the robustness of our watermark detection method against three types of attacks (Table 4). First, we use GPT-3.5-turbo to rewrite the text segments containing the watermark as the paraphrasing attack. The other two attacks randomly swap or delete words at a ratio of 0.2. As expected, rewriting by ChatGPT is the most damaging attack, leading to a decline in detection performance. However, our detection method still significantly outperforms the baseline direct detection across most attack types in terms of TPR. For watermark localization, measured by intersection over union (IoU), our method still generates satisfactory results under these attacks. Overall, the results demonstrate the robustness of our watermark detection approach against various perturbations to the watermarked text.
@@ 5 Conclusion
Appendix A Appendix
A.1 FPR Calibration Function $F$

### v2

@@ Abstract
Text watermarks in large language models (LLMs) are increasingly used to detect synthetic text, mitigating misuse cases like fake news and academic dishonesty. While existing watermarking detection techniques primarily focus on classifying entire documents as watermarked or not, they often neglect the common scenario of identifying individual watermark segments within longer, mixed-source documents. Drawing inspiration from plagiarism detection systems, we propose two novel methods for partial watermark detection. First, we develop a geometry cover detection framework aimed at determining whether there is a watermark segment in long text. Second, we introduce an adaptive online learning algorithm to pinpoint the precise location of watermark segments within the text. Evaluated on three popular watermarking techniques (KGW-Watermark, Unigram-Watermark, and Gumbel-Watermark), our approach achieves high accuracy, significantly outperforming baseline methods. Moreover, our framework is adaptable to other watermarking techniques, offering new insights for precise watermark detection. Our code is publicly available at https://github.com/XuandongZhao/llm-watermark-location.
Efficiently Identifying Watermarked Segments in Mixed-Source Texts
Xuandong Zhaothanks: Co-first authors. UC Berkeley xuandongzhao@berkeley.edu                        Chenwen Liao11footnotemark: 1 Zhejiang University liaochenwen@zju.edu.cn
Yu-Xiang Wangthanks: Co-supervised project. UC San Diego yuxiangw@ucsd.edu                        Lei Li22footnotemark: 2 Carnegie Mellon University leili@cs.cmu.edu
@@ 1 Introduction
However, identifying AI-generated text is becoming increasingly difficult as LLMs reach human-like proficiency in various tasks. One line of research (OpenAI, 2023; Tian, 2023; Mitchell et al., 2023) trains machine learning models as AI detectors by collecting datasets consisting of both human and LLM-generated texts. Unfortunately, these approaches are often fragile (Shi et al., 2024) and error-prone (Liang et al., 2023), ultimately leading OpenAI to terminate its deployed detector (Kelly, 2023). Watermarking has emerged as a promising solution to this challenge [^1] Watermark has been used in industry, e.g. SynthID Dathathri et al. (2024) for Google Gemini.. By embedding identifiable patterns within the generated text, watermarks can signal whether a piece of text originates from an LLM (Dathathri et al., 2024).
Existing watermark detection methods (Aaronson, 2023; Kirchenbauer et al., 2023; Zhao et al., 2023; Kuditipudi et al., 2023; Christ et al., 2023; Hu et al., 2024; Dathathri et al., 2024) are primarily designed for text-level classification, labeling a piece of text as either watermarked or not. However, these methods are insufficient for many real-world scenarios where documents contain mixed-source texts, and only specific sections are LLM-generated. For instance, malicious actors might use LLMs to manipulate certain sections of a news article to spread misinformation. Detecting watermarks within long, mixed-source texts presents a significant challenge, especially when aiming for subsequence-level detection with uncertainty quantification, similar to plagiarism detection systems like “Turnitin [^2] https://www.turnitin.com”. This is because the watermarked signal may be weakened throughout the increasing text length and may not be easily identifiable using conventional detection methods.
@@ 2.1 LLM Watermark and Detection
Red-Green (statistical) watermarking methods partition the vocabulary into two sets, “green” and “red”, using a pseudorandom function $R(h,{\sf k},\gamma)$. This function takes as input the length of the preceding token sequence ($h$), a secret watermark key (${\sf k}$), and the target proportion of green tokens ($\gamma$). During text generation, the logits of green tokens are subtly increased by a small value $\delta$, resulting in a higher proportion of green tokens in the watermarked text compared to non-watermarked text. Two prominent Red-Green watermarking methods are KGW-Watermark (Kirchenbauer et al., 2023, 2024) and Unigram-Watermark (Zhao et al., 2023). KGW-Watermark utilizes $h\geq 1$, considering the prefix for hashing. Unigram-Watermark employs fixed green and red lists, disregarding previous tokens by setting $h=0$ to enhance robustness. Watermark detection in both methods involves identifying each token’s membership in the green or red list
@@ 3 Method
Identifying watermark segments within a long text sequence $y$ presents two key challenges. First, we need to design a classification rule $\mathcal{M}(y)\rightarrow\{0,1\}$ that determines whether $y$ contains a watermark segment. To address this, we propose the Geometric Cover Detector (GCD), which enables multi-scale watermark detection. Second, accurately locating the watermark segments $y_{s_{i}:e_{i}}$ within the full sequence $y$ requires finding the start and end token indices, $s_{i}$ and $e_{i}$, for each watermark segment. We introduce the Adaptive Online Locator (AOL) with the Aligator algorithm to precisely identify the position of the watermarked text span within the longer sequence.
Figure 1: Illustration of the watermark segment detection process. The input sequence could be a mixed-source of watermarked text and unwatermarked text. We use geometric covers to partition the text and detect watermarks in intervals. We also formulate localization as an online denoising problem to reduce computational complexity. The example shown is drawn from the abstract of Bengio et al. (2024), with the watermarked part generated by a watermarked Mistral-7B model.
Algorithm 1 Geometry Cover WM Detection
Input: Text $y$ of length $n$, target false positive rate (FPR) $\tau$, watermark detector score function $\operatorname{Score}$, FPR calibration function $F$
@@ 3.1 Watermark Segment Classification
$\mathcal{I}=\bigcup_{k\in\mathbb{N}\cup{0}}\mathcal{I}^{(k)},$ where $\forall k\in\mathbb{N}\cup{0}$, and
@@
Algorithm 1 outlines our approach. For each segment $\mathcal{I}_{t}:y_{i_{t}:j_{t}}$ in the GC, we first compute a detection score using the appropriate watermark detector for the scheme employed (e.g., Equation 1 for Red-Green Watermark or Equation 2 for Gumbel Watermark). This score, along with the segment itself, is then passed to an FPR calibration function $F$. This function estimates the FPR associated with the segment. Further details on FPR calibration can be found in the Appendix B.
@@ 3.2 Precise Watermark Position Localization
Input: Text $y$, threshold $\zeta$, iterations $m$
@@
Furthermore, relying solely on individual token scores for localization is unreliable due to the inherent noise in the watermarking process. To address this issue, we propose to formulate it as a sequence denoising problem (a.k.a., smoothing or nonparametric regression) so we can provide a pointwise estimate of the expected detection score for each token. Specifically, the denoising algorithm takes a sequence of noisy observations $s_{1},...,s_{n}$ and outputs $\{\theta_{t}\}_{t\in[n]}$ as an estimate to $\{\mathop{\mathbb{E}}[s_{t}]\}_{t\in[n]}$.
@@
$\frac{1}{n}\sum_{t}(\theta_{t}-\mathbb{E}[s_{t}])^{2}=\tilde{O}\Bigg{(}\min\Bigg{\{}n^{-1}\big{(}1+\sum_{t=2}^{n}\mathbf{1}_{\mathbb{E}[s_{t}]\neq\mathbb{E}[s_{t-1}]}\big{)},\\ n^{-1}\vee n^{-2/3}\big{(}\sum_{t=2}^{n}\big{|}\mathbb{E}[s_{t}]-\mathbb{E}[s_{t-1}]\big{|}\big{)}\Bigg{\}}\Bigg{)}.$ (4)
@@
$\frac{1}{j-i}\sum_{t=i}^{j}(\theta_{t}-\frac{1}{j-i}\sum_{t^{\prime}=i}^{j}\mathop{\mathbb{E}}[s_{t^{\prime}}])^{2}\leq\tilde{O}(1/\sqrt{j-i}).$
@@ 4 Experiment
4.1 Experiment Setting
We conduct experiments on three state-of-the-art watermarking methods: Gumbel-Watermark (Aaronson, 2023), KGW-Watermark (Kirchenbauer et al., 2023), and Unigram-Watermark (Zhao et al., 2023) (with $\gamma=0.5$, $\delta=2.0$, and temperature = 1.0). We evaluate our methods on mixed-source texts constructed from the C4 (news-like subset) (Raffel et al., 2020) and Arxiv datasets (Cohan et al., 2018). These texts contain both human-written and LLM-generated (watermarked) segments. Watermarked segments are generated using LLaMA-7B (Touvron et al., 2023) and Mistral (Jiang et al., 2023), and are embedded within longer contexts (10% watermarked ratio to simulate realistic watermark integration), with their positions randomized. For segment detection, we compare our method (GCD) against each watermarking method’s original detector (Vanilla). For localization, we compare AOL with RoBERTa-based classifiers, which are trained to label text segments as watermarked or not based on token watermark scores. We also included WinMax (Kirchenbauer et al., 2024) with window sizes of 1 and 100 (WinMax-1, WinMax-100) as a brute-force baseline ($\tilde{\mathcal{O}}(n^{2})$) for segment detection and localization. We evaluate using True Positive Rate (TPR) at calibrated False Positive Rates (FPRs, both per-segment and document-level), Intersection over Union (IoU) for localization accuracy. We also benchmark detection speed, reporting the average runtime per sample. Further details, including dataset specifics, baseline training, and FPR calibration, are in the Appendix A.
4.2 Detection Results
Watermark Segment Classification Results.
Precise Watermark Position Localization Results.
For the watermark position localization task, we evaluate our proposed method AOL against the baseline method RoBERTa (Table 2). We calculate the average IoU score to quantify the precision of the watermark localization. Our method consistently outperforms the baseline across all test settings. For example, on the C4 dataset using the Mistral-7B model, AOL achieves a substantially higher IoU score of 0.809 compared to 0.301 for RoBERTa. We also test AOL’s ability to detect multiple watermarks by inserting 3x300-token Gumbel watermarks (generated by Mistral-7B) into 6000-token texts. Across 200 samples, the average IoU for detecting the watermarks is 0.802, demonstrating AOL’s effectiveness for multiple watermark detection. Figure 2 provides a case example illustrating the improved localization performance of AOL on the Gumbel watermark with the Mistral-7B model. The upper image shows the boundary effects of using online learning. The lower image demonstrates more precise localization resulting from the circular starting strategy with 10 random starting points. We explored other values for $m$, and our experiments indicate that $m=10$ provides a robust balance between accuracy and computational efficiency for our tested datasets and models.
@@ 4.2 Detection Results
4.3 Detection Efficiency
Method Time (s) TPR IoU
WinMax-1 3643.61 0.99 0.980
WinMax-100 33.90 0.99 0.791
GCD 1.24 0.99 0.387
AOL 2.18 0.99 0.718
Table 3: Runtime and performance comparison of different methods.
Efficiently identifying watermarked segments is a critical goal of our approach. To assess the efficiency of our method, we compare against the brute-force WinMax algorithm Kirchenbauer et al. (2024), a representative sliding window method, with window sizes of 1 and 100 tokens (WinMax-1, WinMax-100). Table 3 presents the results of our comparative evaluation.
The results demonstrate that WinMax achieves high true positive rate (TPR) and intersection over union (IoU) but at a prohibitive computational cost. WinMax-1 requires 3643.61 seconds per sample, while WinMax-100 reduces this to 33.9 seconds. In contrast, our approach (AOL) achieves comparable TPR and a strong IoU (0.718) in just 2.18 seconds, offering a dramatic improvement in computational efficiency. Additionally, we analyze the role of GCD in this efficiency gain. While GCD alone detects watermark presence with an IoU of 0.387, integrating it with AOL significantly enhances localization precision (IoU = 0.718). This ablation study underscores the importance of AOL in refining detection beyond the rough localization provided by GCD.
@@ 4.3 Detection Efficiency
Table 4: Vanilla and GCD watermark segment classification results using the Unigram Watermark on Mistral-7B for different segment-level false positive rate targets, achieved by adjusting score thresholds.
4.4 Detection Results with Different Lengths
@@ 4.5 Detection Robustness Against Attacks
Table 5: Watermark segment classification and localization performance with different attacks, evaluated at three distinct FPR levels (FPR-1/2/3) and IoU scores.
We evaluate the robustness of our watermark detection method against three types of attacks (Table 5). First, we use GPT-3.5-turbo to rewrite the text segments containing the watermark as the paraphrasing attack. The other two attacks randomly swap or delete words at a ratio of 0.2. As expected, rewriting by ChatGPT is the most damaging attack, leading to a decline in detection performance. However, our detection method still significantly outperforms the baseline direct detection across most attack types in terms of TPR. For watermark localization, measured by IoU, our method still generates satisfactory results under these attacks. Overall, the results demonstrate the robustness of our watermark detection approach against various perturbations to the watermarked text.
@@ 5 Conclusion
Limitations
Our methods, GCD and AOL, can be applied to other watermarking schemes as long as they have token-wise detection scores for the sequence, such as Hu et al. (2024) and Zhao et al. (2024). The detection results are constrained by the strength of the original watermark generation and the quality of the prompt text. In some cases, low-quality text produced by the watermark generation method cannot be directly detected using the original detection method. Additionally, positive samples created by inserting the generated watermark paragraph into natural text may not be detectable with our approach. However, these limitations arise from the current limitations of watermark generation and detection methods themselves, which is outside the scope of detecting small watermarked segments within long text, the focus of this work. Therefore, we assume that our method needs only to detect reasonably high quality watermarked text segments embedded in long text.
Appendix A Detailed Experiment Settings
A.1 Datasets and Mixed-source Texts
We utilize two text datasets: C4 (Raffel et al., 2020) and Arxiv (Cohan et al., 2018). The “Colossal Clean Crawled Corpus” (C4) dataset is a large collection of English web text from the public Common Crawl, providing unwatermarked human-written content. We use random samples from its news-like subset. The Arxiv dataset, sourced from arXiv.org and PubMed.com, consists of scientific paper abstracts and articles. Both datasets provide unwatermarked (negative) and partially watermarked (positive) samples. To transform unwatermarked samples into partially watermarked samples, we randomly select 3-5 sentences in a long text and set them as prompts. Then, we generate watermarked text conditioned on the prompts using large language models. The generated responses replace the original suffix sentences after the prompt. To simulate realistic watermark embedding within longer documents and ensure a detectable watermark length, we embed 300-token watermarks within 3000-token contexts (10% ratio). This ratio is designed to represent a plausible level of watermark integration in texts such as articles or web pages, while simultaneously providing a sufficient length for reliable detection. The watermark was randomly positioned within the longer context, and these positions were recorded for subsequent localization evaluation. Our objective is to determine the presence of watermarked text within a document and accurately locate its segment. For testing, we utilized 500 samples per dataset.
A.2 Language Models and Watermarking Methods
We use the publicly available LLaMA-7B (Touvron et al., 2023) and Mistral (Jiang et al., 2023) models. To verify the general applicability of the watermark detection methods, we select three watermarking techniques: Gumbel-Watermark (Aaronson, 2023), KGW-Watermark (Kirchenbauer et al., 2023), and Unigram-Watermark (Zhao et al., 2023). These methods represent the state-of-the-art watermarking approaches for LLMs, offering high quality, detectability, and robustness against adversarial attacks. For all watermarking generations, we configure the temperature to 1.0 for multinomial sampling. Additionally, for KGW-Watermark and Unigram-Watermark, we set the green token ratio $\gamma$ to 0.5 and the perturbation $\delta$ to 2.0.
A.3 Baselines
In watermark segment detection, we use the original watermark detector in each watermarking method as the Vanilla baseline to compare with our approach GCD. In watermark segment localization, we use RoBERTa (Liu et al., 2019) models for comparing with our method AOL. We train each RoBERTa (designed for different watermarking methods) to predict whether a sequence is a watermarked sequence or not, given the watermark detection scores $r$ for each token. We add an extra fully connected layer after getting the representation of the [CLS] token. We construct 1000 training samples with 60 token scores as input and the binary label of this segment as the label. We train the RoBERTa model for 20 epochs and enable early stopping if the loss converges. It can reach over 90% accuracy in the training set. During testing on mixed-source text, we employ the sliding window idea to test each chunk for watermarks and then calculate the IoU score. We also conducted experiments with the WinMax algorithm (Kirchenbauer et al., 2024), a representative brute-force method with time complexity of $\tilde{\mathcal{O}}(n^{2})$). We tested two window sizes: 1 token (WinMax-1) and 100 tokens (WinMax-100).
A.4 Evaluation
For the watermarked text classification task, we report the true positive rates (TPR) based on different specified false positive rates (FPR). Maintaining a low FPR is critical to ensure that human-written text is rarely misclassified as LLM-generated text. Since the FPR at the per-instance level differs from the document-level FPR, we calibrate FPR to three distinct levels in each scenario to enable fair comparisons. Specifically, we manipulate the pre-segment FPR (Seg-FPR) by adjusting the threshold parameter $\tau$ as outlined in Algorithm 1. Then, we can get the empirical document FPR (Doc-FPR) by evaluating our method GCD based on pure natural text. For Vanilla, we set the FPR according to GCD’s empirical FPR and subsequently test for its empirical TPR. For locating specific watermark segments, we calculate the Intersection over Union (IoU) score to measure the accuracy of watermark segment localization. The IoU score computes the ratio of the intersection and union between the ground truth and inference, serving as one of the main metrics for evaluating the accuracy of object detection algorithms:
$\operatorname{IoU}$ $=\frac{\text{Area of Intersection}}{\text{Area of Union}}$
$=\frac{|\text{Detected Tokens}\cap\text{Watermarked Tokens}|}{|\text{Detected Tokens}\cup\text{Watermarked Tokens}|}$
Finally, to assess inference speed, we calculate the average time to detect a single sample on a server equipped with an AMD EPYC 9354 32-Core Processor and Nvidia A6000 GPUs.
Appendix B FPR Calibration Function $F$